# FedCLEAN: byzantine defense by CLustering Errors of Activation maps in Non-IID federated learning environments

### v1

@@ 3.1.1 Adversary’s Objective
In this work, we consider a "Byzantine" adversary $A$, who wishes to carry out Byzantine attack during the training phase of an FL model. To do so, $A$ controls a number of the FL clients who submit poisoned model updates to $S$ during the FL process with the objective of disturbing the convergence of the global model. Formally, $A$ attempts to fulfill the following objectives:
  • O1: If the attack succeeds, the global model’s accuracy on test data should collapse.
@@ 3.1.2 Adversary’s Capabilities
As in [3, 46, 22, 11, 24, 10], we consider a strong adversary who has full control over $m<N/2$malicious" clients. $A$ has full access to these clients’ datasets, models, and communication channels. In particular, $A$ can conduct training using local datasets when clients are selected by the server, freely modify the weights of the resulting model updates, and send back modified updates to the server $S$ for aggregation. In accordance with Shannon’s Maxim, $A$ also has full knowledge of the aggregation method used by $S$ and of the defense this one deploys but cannot access or alter its parameters. They also have no access to other benign clients’ updates or datasets.
@@ 3.2 Security Requirements
To effectively defend against a Byzantine attacker, it is required that a filtering-based defense system $D$ satisfies the following requirements [46]:
  • R1: Block all malicious clients. As discussed above, the Byzantine adversary $A$ is able to break the model given one single successful round of attack. As a result, in order to be robust to such an adversary, $D$ must detect all malicious clients updates at any round.
  • R2: Do not disrupt training. $D$ should not negatively impact the training. In other words, $D$ must have high precision and not misclassify benign clients. It must also preserve the test accuracy achieved in the no-attack and no-defense scenario.
We will see that our proposal responds to all these objectives and security requirements.
4 Proposed defense: FedCLEAN
4.1 FedCLEAN Principle
The idea behind FedCLEAN is to analyze the normalized activation maps (NAMs), obtained by subtracting the GeoMed value from activation maps of one or several layers of $M_{G}$ for some given inputs referred as a trigger set $T_{s}=(X_{T},Y_{T})$, where $X_{T}$ and $Y_{T}$ are the set of samples and their labels, respectively. We assume that the NAMs obtained with malicious model updates will correspond to outliers that can be identified by means of an CVAE and GeoMed based anomaly detection strategy.
More clearly, before the FL process begins : FedCLEAN : i) makes a copy of the initial global model; ii) trains the copy locally for several epochs on a small dataset called the "trigger set"; iii) extracts the NAMs associated with the trigger set from different epochs of training; iv) trains a CVAE to reconstruct these NAMs, using their corresponding labels $Y_{T}$ as a condition. The training process is detailed in 4.2. By doing so, the CVAE learns the latent distribution of NAMs associated with each class.
Once trained, the CVAE is used by $S$ at each epoch to reconstruct for each client updates the NAMs corresponding to the trigger set. Malicious NAMs, which fall out of the learned distribution, result in high reconstruction errors. To discriminate malicious NAMs from benign ones, FedCLEAN makes use of a clustering procedure relying on an ad-hoc trust propagation algorithm. We discuss it in subsection 4.3.
To provide a formal description of FedCLEAN, let us consider a classification task and $N$ clients $\{C_{k}\}_{k=1..N}$. FedCLEAN operates according to the following steps :
@@ 4.1 FedCLEAN Principle
    FL and FedCLEAN Initialization
@@
      $S$ initializes the global FL model $M_{G}^{0}$ according to a common process. By assigning, for instance, random values to its weights.
@@
      $S$ then trains the CVAE. To do so, it first extracts the activation maps $A_{S}$ of one or more layers from $M_{G}^{{}^{\prime}0}$ feeding it with the trigger set $T_{s}$. Then, it normalizes them using GeoMed such as: $\hat{A}_{S}=\sigma(A_{S}-\text{GeoMed})$ (where $\sigma$ is the sigmoid function). $S$ finally trains the CVAE on $\hat{A}_{S}$.
@@
      $S$ then sends $M_{G}^{0}$ to $U$ randomly selected clients ($U<N$), to initiate the first FL training round.
@@
    FL training process for each round $t>1$
@@
      Secure FedCLEAN selection
@@
        Inspection the activation maps for each client $u$: $S$ extracts the activation maps $A_{u}^{t}$ obtained by feeding the trigger set $T_{s}$ into $M^{t}_{u}$.
@@
        • $S$ computes the activation maps’ GeoMed: $\text{GeoMed}(A_{1}^{t},...,A_{U}^{t})$ and normalizes the activation maps of each client: $\hat{A}_{u}^{t}=\sigma(A_{u}^{t}-\text{GeoMed})$.
        • $S$ feeds its CVAE with $\hat{A}_{u}^{t}$ conditioned on $Y_{t}$ and assigns to each client a reconstruction error score $\epsilon_{u}^{t}=\overline{MSE}(\hat{A}_{u}^{t},\tilde{A}_{u}^{t})$ where $\tilde{A}_{u}^{t}$ is the CVAE reconstruction of $\hat{A}_{u}^{t}$, and $MSE$ the mean square error function.
@@
      Filtering: Given the pairs $(M_{u}^{t},\epsilon_{u}^{t})$, $S$ performs our trust propagation clustering to filter out clients that are considered malicious and discards their updates.
@@
As exposed, the performance of FedCLEAN relies on two main components a CVAE and a trust propagation clustering algorithm used to discriminate malicious from benign updates. We come back to how they work in the following sections.
4.2 CVAE training constraints
As stated above, the CVAE model is trained by the server $S$ considering a simulated a benign model $M_{G}^{{}^{\prime}0}$ based on the server trigger set $T_{s}$. One of the constraints we assumed in building FedCLEAN is that $T_{s}$ is of very small size compared to the client data. To train the CVAE well, one thus needs to augment the training set. The solution we adopted is to use the activation maps (AMs) computed at different rounds of the training of $M_{G}^{{}^{\prime}0}$. But by doing so, we have to take into account the temporal drift of $M_{G}^{{}^{\prime}0}$ activation maps that impact CVAE generalization capabilities. In fact, AMs are more heavily affected by noise and very different at the very beginning of the training of $M_{G}^{{}^{\prime}0}$ than after a certain number of epochs. To overcome this issue, and as illustrated in Fig. 2 the CVAE training set is constituted of activation maps collected after some warmup epochs of the training of $M_{G}^{{}^{\prime}0}$. At each epoch, AMs are GeoMed normalized so as to better control AMs amplitude differences along the training of $M_{G}^{{}^{\prime}0}$. Indeed, with a GeoMed recomputed at each round, we have found that the distribution of the deviations from GeoMed remains stable. Anyway, even with a small trigger set, this technique allows to constitute a training set that captures the distribution of benign NAMs assuming certain hypotheses.
@@ 4.2 CVAE training constraints
A second issue we had to face when training the CVAE is that it is sensitive to "posterior collapse" [29, 9, 14], which causes it to converge to minima that ignore training inputs. More clearly, the traditional training of a CVAE, consists in maximizing the Evidence Lower Bound (ELBO) in the following loss:
@@
where: $g_{\theta}$ is the CVAE decoder; $q_{\phi}$ is the latent space distribution given by the CVAE encoder; $z$ is a latent representation sampled from this distribution; and, $(x,y)$ the CVAE input sample and its label; $D_{KL}$ is the Kullback-Leibler divergence measuring the divergence with a prior distribution $p$, in our case a standard normal distribution. As seen, this loss is constituted of a first term to minimize the CVAE reconstruction error and of a second to enforce a Gaussian structure for the latent space representation. In our context, NAMs give rise to numerically very small MSE values, so the optimizer focuses on minimizing the second term, leading to the so called "posterior collapse" issue. To solve this issue, it has been proposed to add a regularization weight to the divergence term in the loss function [29]. In our work, we adopted a KLD annealing approach with an adaptive regularization weight $\beta^{t}$ like [9]
@@
where, $t$ is the current training epoch. CVAE starts to train with $\beta^{0}=0$ so that it learns to produce meaningful representations of the feature space and make predictions based on such representations. Then,$\beta$ is gradually increased by a fixed increase to enforce the structure of the latent space.
@@ 4.3 Clustering using trust propagation
A commonplace practice in CVAE anomaly detection and FL defenses [6, 24, 37] is to threshold error values to decide of an anomaly. Usually, this is the average reconstruction error of all clients. However because of this, in the absence of an attacker, nearly half of the clients even if they are all benigns are wrongly flagged as malicious. Well-adjusted static thresholds could allow for better filtering. In an IID scenario, such reference values could be implied from the error values of CVAE during training. However, in a Non-IID scenario, real client reconstruction error values cannot be deduced from the sample training scenario. We thus opted for a simple clustering algorithm instead.
@@
To avoid systematically excluding benign clients, our clustering algorithm is able to automatically output either one or two clusters depending on whether malicious updates are present. This is a challenge as deciding the number of clusters based on data properties is an active problem in cluster analysis [21], and most algorithms require fixing the number of clusters beforehand as a user-defined parameter [21]. In addition, clustering algorithms vary widely in terms of "parameter reliance", i.e. the dependence on user-defined parameters and the tolerance to different ranges of such parameters. Our proposal has a single parameter that is interpretable (i.e. not rely on guessing).
@@
It sequentially retrieves all benign clients participating in training, through what we call a "trust propagation" algorithm. It relies on a single relative parameter, which is a fraction in $[0,1]$ that can be selected in a data-agnostic manner, which it then uses to dynamically compute the distance measures it needs to perform clustering. Our idea relies on the fact that when reconstructing benign client activation maps, the measured reconstruction error is a residual from the randomness and uncertainty within CVAE. More clearly, when reconstructing out-of-distribution attacker NAMs, the reconstructed vector is an arbitrary response. Thus, the reconstruction error we measure is akin to measuring error to a random vector. As a result, assuming we are working with an autoencoder that has properly fitted benign NAMs distributions, we expect the error from the reconstruction of benign NAMs to be of a lower order of magnitude than those from anomalous NAMs. In other words: in the reconstruction error space, there is an identifiable relative error gap between the two.
@@
Our algorithm works as follows. It sorts the client reconstruction errors, and selects the lowest reconstruction error client as the core of the benign client group. Then, in the reconstruction error space, if a client is within a certain distance $\delta$ of the last accepted client, he is added to the group. The iterations stop when the next client is at $d>\delta$ distance from the current client. We assume in that case that the gap between error values corresponds to that between attackers and benigns, and block all remaining clients. $\delta$ is computed dynamically as a proportion of the size of the error space for the current round, as $\delta=\lambda(max(\epsilon_{u}^{t})-min(\epsilon_{u}^{t}))_{0<u<U}$, where $\lambda\in[0,1]$ is a hyper-parameter fixed at the start of the defense system. The full pseudo-code for the trust propagation algorithm is detailed in Alg. 1.
@@
$\delta$, $clients\_sorted$, $sorted\_re$
$benign\_clients\leftarrow clients\_sorted[0]$
$i\leftarrow 0$
$i_{max}\leftarrow\text{len}(clients\_sorted)-1$
while ($i\neq i_{max}$) and ($sorted\_re[i+1]<sorted\_re[i]+\delta$do
     $benign\_clients.\text{append}(clients\_sorted[i+1])$
     $i\leftarrow i+1$
@@
return $benign\_clients$
@@ 5.1.1 FL setting
We evaluate our proposition on Non-IID image classification tasks using two public datasets : MNIST [17] and FashionMNIST [51] datasets. The FL is orchestrated by a server that applies FedCLEAN as defense $D$ before each aggregation step with 100 total clients. 50% of clients are randomly selected to participate in training at each round, and we use FedAVG [41] to aggregate model updates. As global model $M_{G}$, we use a standard CNN architecture that achieves good results on both datasets in the absence of attackers. In all our experiments, we set 30% of the client population to be attackers.

### v2

@@ 3.1.1 Adversary’s Objective
In this work, we consider a "Byzantine/backdoor" adversary $A$, who wishes to carry out Byzantine/backdoor attack during the training phase of an FL model. To do so, $A$ controls a number of the FL clients who submit poisoned model updates to $S$ during the FL process with the objective of disturbing the convergence of the global model. Formally, for Byzantine attacks, $A$ attempts to fulfill the following objectives:
  • O1: If the attack succeeds, the global model should have bad accuracy on test data.
@@
And, for a backdoor to be successful :
  • O3: Achieve a good attacker success rate (ASR), which corresponds to the ratio of test samples with the backdoor trigger that the model successfully classifies in the target class.
@@ 3.1.2 Adversary’s Capabilities
As in [54, 15, 55, 13, 12, 14], we consider a strong adversary who has full control over $m<N/2$ "malicious" clients; $N$ being the total number of clients. $A$ has full access to datasets, models, and communication channels of these clients. In particular, $A$ can conduct training using local datasets when clients are selected by the server, freely modify the weights of the resulting model updates, and send back modified updates to the server $S$ for aggregation. In accordance with Shannon’s Maxim, $A$ also has full knowledge of the aggregation method used by $S$ and of the defense this one deploys but cannot access or alter its parameters. They also have no access to other benign clients’ updates or datasets.
@@ 3.2 Security Requirements
To effectively defend against a Byzantine or backdoor attacker, it is required that a filtering-based defense system $D$ satisfies the following requirements [15]:
  • R1: Block all malicious byzantine clients. As discussed above, the Byzantine adversary is able to break the model given one single successful round of attack. As a result, in order to be robust to such an adversary, $D$ must detect all of these malicious client updates at any round.
  • R2: Do not disrupt training. $D$ should not negatively impact the training. In other words, $D$ must have high precision and not misclassify benign clients. It must also preserve the test accuracy achieved in the baseline no-attack and no-defense scenario.
  • R3: Lower backdoor attacker success rate. In the presence of the defense, the backdoor accuracy measured on the test set should be significantly lower.
We will see in the sequel that our proposal fulfills these three security requirements.
4 Proposed defense: FL-CLEANER
4.1 FL-CLEANER Basic Principles
The idea behind FL-CLEANER is to analyze the normalized activation maps (NAMs), obtained by subtracting the GeoMed value from activation maps of one or several layers of $M_{G}$ for some given inputs referred as a trigger set $T_{s}=(X_{T},Y_{T})$, where $X_{T}$ and $Y_{T}$ are the set of samples and their labels, respectively. We assume that NAMs obtained with malicious model updates will correspond to outliers that can be identified by means of a CVAE and GeoMed based anomaly detection strategy.
In a simplified view, before the FL process begins, FL-CLEANER :
@@ 4.1 FL-CLEANER Basic Principles
    Makes a copy of the initial global model.
  2. 2.
    Trains this copy locally for several epochs on a "trigger set" that is a very small dataset.
  3. 3.
    Extracts NAMs associated with the trigger set from different epochs of training
  4. 4.
    Trains a CVAE to reconstruct these NAMs, using their corresponding labels $Y_{T}$ as a condition. The training process is detailed in 4.2.
By doing so, the CVAE learns the latent distribution of NAMs associated with each class.
Once trained, $S$ uses CVAE at each epoch to reconstruct NAMs corresponding to the trigger set for each client updated model $S$ received. Malicious client NAMs, which fall out of the learned distribution, result in high reconstruction errors. To discriminate malicious NAMs from benign ones, FL-CLEANER makes use of a clustering procedure relying on an ad-hoc trust propagation algorithm. We present and discuss it in subsection 4.3.
To go in further details and provide a formal description of FL-CLEANER, let us consider a FL scenario with a classification task and $N$ clients $\{C_{k}\}_{k=1..N}$. FL-CLEANER operates according to the following steps:
  1. 1.
    FL and FL-CLEANER Initialization ($t=0$)
@@
      $S$ initializes the global FL model $M_{G}^{0}$ in a classic way assigning, for instance, random values to its weights.
@@
      $S$ then trains a CVAE model. To do so, it first extracts the activation maps $A_{S}$ of one or more layers from $M_{G}^{{}^{\prime}0}$ feeding it with the trigger set $T_{s}$. Then, it normalizes them using GeoMed such as: $\hat{A}_{S}=\sigma(A_{S}-\text{GeoMed})$, with $\sigma$ is the sigmoid function. $S$ finally trains CVAE on $\hat{A}_{S}$.
@@
      $S$ initiates the first FL training round by sending $M_{G}^{0}$ to $U$ randomly selected clients ($U<N$),.
@@
    Round $t\geq 1$ of the FL training process
@@
      Secure FL-CLEANER selection
@@
        Activation map inspection for each client $u$ : $S$ extracts the activation maps $A_{u}^{t}$ obtained by feeding the trigger set $T_{s}$ into $M^{t}_{u}$.
@@
        • $S$ computes $\text{GeoMed}(A_{1}^{t},...,A_{U}^{t})$, that is the GeoMed centroid of all activation maps. Then it normalizes the activation maps of each client: $\hat{A}_{u}^{t}=\sigma(A_{u}^{t}-\text{GeoMed})$.
        • $S$ feeds its CVAE with $\hat{A}_{u}^{t}$ conditioned on $Y_{t}$ , the label, and assigns to each client a reconstruction error score $\epsilon_{u}^{t}=\overline{MSE}(\hat{A}_{u}^{t},\tilde{A}_{u}^{t})$ where $\tilde{A}_{u}^{t}$ is the CVAE reconstruction of $\hat{A}_{u}^{t}$, and $MSE$ the mean square error function.
@@
      Filtering: Given the pairs $(M_{u}^{t},\epsilon_{u}^{t})$, $S$ performs our trust propagation clustering (see next Section) to filter out clients that are considered malicious and discards their updates.
@@
As exposed, the performance of FL-CLEANER to discriminate malicious from benign updates rely on two main components: a CVAE model and a trust propagation clustering algorithm. We come back to how they work in the following sections.
4.2 CVAE training
As stated above, the CVAE model is trained by the server $S$ considering a simulated benign model $M_{G}^{{}^{\prime}0}$ based on the server trigger set $T_{s}$. In this work we assume that $T_{s}$ is of very small size compared to the client data. This ensures the real-life applicability of our solution as servers may not have access to the large datasets other methods in literature require. Moreover, servers may have some data of their own or access to trusted clients or can form such datasets from open-source data. The influence of the trigger set size and content is not included in our study due to space limitations. However, our experiments showed that it is possible to achieve good defense results against Byzantine attackers with as little as 250 images randomly sampled from the server test set and 500 images for backdoor attacks.
To augment the number of training samples available to the CVAE with a small dataset, the solution we adopted is to use the activation maps (AMs) computed at different rounds of the training of $M_{G}^{{}^{\prime}0}$. AMs are more heavily affected by noise and vary a lot from one epoch to another at the very beginning of the training of $M_{G}^{{}^{\prime}0}$. To account for this, as whown in in Fig. 3, the CVAE training set is constituted of activation maps collected after some warmup epochs of the training of $M_{G}^{{}^{\prime}0}$. Additionnally, AMs are GeoMed normalized into NAMS at each epoch so as to better control AMs amplitude differences along the training of $M_{G}^{{}^{\prime}0}$. Doing so also helps to take into account the temporal drift of $M_{G}^{{}^{\prime}0}$ activation maps. Indeed, the distribution of the deviations from GeoMed recomputed at each iteration remains stable throught the epochs. That is not the case of the distribution of the AMs which drifts over time. We will see that even with a small trigger set, these technique allows to constitute a training set that well captures the distribution of benign NAMs, and ensures good CVAE generalization capabilities.
@@ 4.2 CVAE training
A second issue to consider is that CVAE training is sensitive to "posterior collapse" [56, 57, 58], causing it to converge to minima that ignore training inputs. Indeed, the traditional training of a CVAE, consists in maximizing the Evidence Lower Bound (ELBO) in the following loss:
@@
where: $g_{\theta}$ is the CVAE decoder; $q_{\phi}$ is the latent space distribution given by the CVAE encoder; $z$ is a latent representation sampled from this distribution; and, $(x,y)$ the CVAE input sample and its label; $D_{KL}$ is the Kullback-Leibler divergence measuring the divergence with a standard normal prior distribution $p$. As exposed, the first loss term minimizes the CVAE reconstruction error and the second enforces a Gaussian structure for the latent space representation. In our context, NAMs give rise to numerically very small reconstruction error values making the optimizer to focus on minimizing the second term: the so called "posterior collapse" issue. To solve this issue, we add a regularization weight to the divergence term in the loss function [56]. In our work, we adopted a KLD annealing approach with an adaptive regularization weight $\beta^{t}$ as in [57]
@@
where, $t$ is the current training epoch. CVAE starts its training with $\beta^{0}=0$ Then, $\beta$ is gradually increased by a fixed increase to enforce the structure of the latent space.
@@ 4.3 Clustering using trust propagation
A commonplace practice in CVAE anomaly detection and FL defenses [16, 12, 23] is to threshold error values to decide of an anomaly. Usually, this is the average reconstruction error of all clients [16, 12]. A result of this is that in the absence of an attacker almost half of the clients are wrongly flagged as malicious, even if they are benign. Well-adjusted static thresholds could allow for better filtering. In an IID scenario, such reference values could be implied from the error values of CVAE during training. However, in a Non-IID scenario, real client reconstruction error values cannot be deduced from the sample training scenario. We opted for a simple clustering algorithm instead.
@@
Our clustering algorithm automatically outputs either one or two clusters depending on whether malicious updates are present. Notice that deciding the number of clusters based on data properties is an active problem in cluster analysis [59]. Indeed, most algorithms require fixing the number of clusters beforehand as a user-defined parameter [59] with clustering performance that varies widely in terms of "parameter reliance", i.e. parameter tolerance ranges. Our proposal has a single parameter that is interpretable (i.e. not rely on guessing) and fixes the number of clusters automatically.
@@
The basic idea of our solution relies on the fact that for benign client’s activation maps the measured reconstruction error is a residual from the randomness and uncertainty within CVAE and that the reconstructed vector in the case of an out-of-distribution attacker NAMs is an arbitrary response akin to measuring error to a random vector. Assuming that the CVAE has properly fitted benign NAMs distribution, expected reconstruction errors of benign NAMs are of a lower order of magnitude than for anomalous NAMs. In other words: in the reconstruction error space, there is an identifiable relative error gap between the two, as exemplified in Figure 4 in the case of a sign flipping attack (sign inversion of mdoel weights).
@@
Based on this, we propose a "trust propagation" algorithm the principle of which consists in sequentially retrieving all benign clients participating in training. First, it sorts the client reconstruction errors, and selects as the core of the benign client cluster the one with the lowest error score. Then, in the reconstruction error space, if a client is within a certain distance $\delta$ of the last accepted client, it is added to the group. The iterations stop when the distance $d$ of the next client from the current client is such as $d>\delta$. This gap between error values corresponds to the frontier between attackers and benigns. All the remaining clients constitute the attacker cluster, and are blocked. Our algorithm depends on one single parameter $\delta$ which can be dynamically computed as a proportion of the size of the error space for the current round, such as:
$\delta=\lambda(max(\epsilon_{u}^{t})-min(\epsilon_{u}^{t}))_{0<u<U}$ (3)
Where $\lambda\in[0,1]$ is a hyper-parameter fixed in a data-agnostic manner at the start of the defense system. The full pseudo-code for the trust propagation algorithm is detailed in Alg. 1.
Figure 4: Histogram of reconstruction error values measured for benign clients and attackers performing a Sign Flipping attack on MNIST, considering a trigger set $T_{s}$ of 250 samples.
By sequentially building the trusted clients cluster rather than fixing a threshold, our detection handles attacks but also normal scenarios where the highest error clients are still benign. We observed that benign clients are very rarely misclassified with our method, excluding less than 1% of benign clients in the worst case (see 5.4).
@@
[…]
$\text{benign\_clients}\leftarrow\text{clients\_sorted[0]}$
$\text{i}\leftarrow\text{0}$
$\text{i\_{max}}\leftarrow\text{len(clients\_sorted)-1}$
[…]
  benign_clients.append(clients_sorted[i+1])
  $\text{i}\leftarrow\text{i + 1}$
@@
return  benign_clients
@@ 5.1.1 FL setting
We evaluated FL-CLEANER on Non-IID image classification tasks using two public datasets : MNIST [60] and FashionMNIST [61]. In the following, FL is orchestrated by a server that applies FL-CLEANER as defense $D$ before each aggregation step with 100 total clients. 50% of clients are randomly selected to participate in training at each round, and we use FedAVG [8] to aggregate model updates. As global model $M_{G}$, we use a standard CNN architecture that achieves good results on both datasets in the absence of attackers. In all our experiments, we set 30% of the client population to be attackers, in accordance with previous works in literature [16].